# Global solutions to cross-diffusion systems with independent advections in one dimension

### 20 Mar 2026 v1

@@ 3.1. A priori estimates
$\|s_{\varepsilon}^{\frac{\gamma+\alpha}{2}+1}\|_{L^{\infty}(\mathbb{R})}\leq\|\partial_{x}s_{\varepsilon}^{\frac{\gamma+\alpha}{2}+1}\|_{L^{1}(\mathbb{R})}\leq\left(\frac{\gamma+\alpha}{2}+1\right)\|s_{\varepsilon}^{\frac{3}{2}}\|_{L^{2}(\mathbb{R})}\,\|s_{\varepsilon}^{\frac{\gamma+\alpha-3}{2}}\,\partial_{x}s_{\varepsilon}\|_{L^{2}(\mathbb{R})},$
@@
$\int_{\mathbb{R}}s_{\varepsilon}^{\frac{\gamma-3\alpha}{2}}\mathop{}\!\mathrm{d}x\leq\left(\int_{\mathbb{R}}s_{\varepsilon}^{\frac{\gamma}{2}}\mathop{}\!\mathrm{d}x\right)^{\beta}\,M^{1-\beta},\qquad\qquad\beta=\frac{\frac{\gamma-3\alpha}{2}-1}{\frac{\gamma}{2}-1},$
@@
We finally observe that $M^{2(1-\beta)}\leq 1+M^{2}$ and $\gamma\geq 1$ and $\beta\geq\frac{4}{5}$ implies $\gamma^{\frac{4}{\beta}}\leq\gamma^{5}$ so that
$\partial_{t}\int_{\mathbb{R}}s_{\varepsilon}^{\gamma}\mathop{}\!\mathrm{d}x\leq\frac{1+M^{2}}{\delta_{1}^{2}}\,\|s_{\varepsilon}^{\frac{3}{2}}\|^{2}_{L^{2}(\mathbb{R})}\,\left(1+\gamma^{5}\,\left(\int_{\mathbb{R}}s_{\varepsilon}^{\frac{\gamma}{2}}\mathop{}\!\mathrm{d}x\right)^{2}\right).$
@@
Step 6. Proof of (E5), (E6). Clearly, $\{u_{\varepsilon}\log u_{\varepsilon}\}$, $\{v_{\varepsilon}\log v_{\varepsilon}\}$ are bounded in $L^{\infty}((0,T)\times\mathbb{R})$ by (E4). Hence, to obtain (E5) we need to prove that $\{u_{\varepsilon}\log u_{\varepsilon}\}$, $\{v_{\varepsilon}\log v_{\varepsilon}\}$ are bounded in $L^{\infty}(0,T;L^{1}(\mathbb{R}))$. We observe that by Lemma 3.3 with $|y|^{-}=-\min(0,y)$
$\int_{\mathbb{R}}u_{\varepsilon}|\log u_{\varepsilon}|\mathop{}\!\mathrm{d}x=\int_{\mathbb{R}}u_{\varepsilon}\log u_{\varepsilon}\mathop{}\!\mathrm{d}x+2\int_{\mathbb{R}}u_{\varepsilon}|\log u_{\varepsilon}|^{-}\mathop{}\!\mathrm{d}x\leq\\ \leq\int_{\mathbb{R}}u_{\varepsilon}\log u_{\varepsilon}\mathop{}\!\mathrm{d}x+\int_{\mathbb{R}}u_{\varepsilon}\,W(x)\mathop{}\!\mathrm{d}x+\frac{2}{e}\int_{\mathbb{R}}e^{-W(x)}\mathop{}\!\mathrm{d}x,$
where $W:\mathbb{R}\to[0,\infty)$ satisfies $e^{-W(x)}\in L^{1}(\mathbb{R})$. The same inequality holds for $v_{\varepsilon}$. Therefore, the claim for $\alpha>\frac{1}{3}$ follows from (3.3) and the bound (E1) which in this case was proved independently (we choose $W(x)=|x|^{2}$) while for $\alpha\leq\frac{1}{3}$ the claim follows from (3.10) (we choose $W(x)=|x|/2$) and the proof of (E5) is concluded.
To prove (E6), we split the integral for three sets as follows
$\int_{\mathbb{R}}u_{\varepsilon}|\log u_{\varepsilon}||x|^{\sigma}\mathop{}\!\mathrm{d}x\leq\int_{u_{\varepsilon}\geq 1}u_{\varepsilon}|\log u_{\varepsilon}||x|^{\sigma}\mathop{}\!\mathrm{d}x+\int_{u_{\varepsilon}\in[e^{-|x|^{\sigma}},1]}u_{\varepsilon}|\log u_{\varepsilon}||x|^{\sigma}\mathop{}\!\mathrm{d}x+\\ +\int_{u_{\varepsilon}\leq e^{-|x|^{\sigma}}}u_{\varepsilon}|\log u_{\varepsilon}||x|^{\sigma}\mathop{}\!\mathrm{d}x=:I_{1}+I_{2}+I_{3},$
where $\sigma=1$ if $\alpha>\frac{1}{3}$ and $\sigma=\frac{1}{2}$ if $\alpha\leq\frac{1}{3}$. For $I_{1}$, we can estimate $|\log u_{\varepsilon}|\leq 1+u_{\varepsilon}\leq 1+\|u_{\varepsilon}\|_{L^{\infty}}$ by (E4) and $|x|^{\sigma}\leq 1+|x|^{2\sigma}$ so we can use (E1) to control the integral. For $I_{2}$, the set of integration implies $|\log u_{\varepsilon}|\leq|x|^{\sigma}$ so we can use again (E1) to control the term. Finally, to bound $I_{3}$, we write $u_{\varepsilon}|\log u_{\varepsilon}||x|^{\sigma}=\sqrt{u_{\varepsilon}}\,\sqrt{u_{\varepsilon}}|\log u_{\varepsilon}|\,|x|^{\sigma}\leq C\,|x|^{\sigma}\,e^{-|x|^{\sigma}/2}$ where $C=\sup_{x\in[0,1]}\sqrt{x}\,|\log x|$. The proof for $v_{\varepsilon}$ is the same.
Step 7. Proof of (E3) and (E8). For (E3), we observe that
@@
Step 8. Proof of (E9). We have $\partial_{t}u_{\varepsilon}=\partial_{x}F_{\varepsilon}$ where
@@
Step 9. Proof of (E10). Since $\alpha\leq 2$ we use that $\partial_{x}s_{\varepsilon}=\frac{2}{\alpha}\,s_{\varepsilon}^{1-\frac{\alpha}{2}}\,\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}$ so $\{\partial_{x}s_{\varepsilon}\}$ is bounded in $L^{2}((0,T)\times\mathbb{R})$ by (E2) and (E4). ∎
@@ 3.2. Basic compactness
    $\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},\!v_{\varepsilon}]\!\to\!\partial_{x}\mathcal{V}^{1}[u,\!v],\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},\!v_{\varepsilon}]\!\to\!\partial_{x}\mathcal{V}^{2}[u,\!v]$ a.e. and strongly in $L^{p}_{\operatorname{loc}}((0,T)\!\times\!\mathbb{R})$,
@@
Concerning the overline notation, we use the it to denote weak limits which exist by the a priori estimates but we do not know their representation. In particular, we do not know that $\overline{|\partial_{x}s^{\alpha}|^{2}}=|\partial_{x}s^{\alpha}|^{2}$ or $\overline{u\,\partial_{x}s^{\alpha}}=u\,\partial_{x}s^{\alpha}$. Before proceeding to the proof, we will state a version of the Aubin-Lions Lemma taken from [45, Corollary 7.9] in the case that the time derivative is a measure.
@@
$\frac{1}{1+\frac{\alpha}{2}}\int_{0}^{T}\int_{\mathbb{R}}\partial_{t}s_{\varepsilon}^{1+\frac{\alpha}{2}}\,\varphi\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}t\leq\|\partial_{t}s_{\varepsilon}\|_{L^{2}(0,T;H^{-1}(\mathbb{R})}\,\|s_{\varepsilon}^{\frac{\alpha}{2}}\|_{L^{2}(0,T;H^{1}(\mathbb{R})}\,\|\varphi\|_{C(0,T;W^{1,\infty}_{0}(B_{R}))}$
@@
$\|\partial_{x}K^{1,1}\ast(u_{\varepsilon}-u)\|_{L^{1}((0,T)\times\Omega)}\leq\\ \leq\|(\partial_{x}K^{1,1}\,\chi_{R})\ast(u_{\varepsilon}-u)\|_{L^{1}((0,T)\times\Omega)}+\|(\partial_{x}K^{1,1}\,(1-\chi_{R}))\ast(u_{\varepsilon}-u)\|_{L^{1}((0,T)\times\Omega)}.$
The first term converges to 0 for each $R>0$ (indeed, by a standard version of the Aubin-Lions lemma, it has a subsequence converging to 0 so we deduce that the whole sequence converges to 0; in order to apply the Aubin-Lions lemma, the spatial regularity is obtained from $\partial_{x}K^{1,1},\partial_{x}^{2}K^{1,1}\in L^{\infty}(\mathbb{R})$ while the temporal regularity is guaranteed by (E9)). The second term is arbitrarily small by the assumption $\|\partial_{x}K^{1,1}\,\mathds{1}_{|x|\geq R}\|_{L^{1}(\mathbb{R})}\to 0$ as $R\to\infty$ (Assumption 1.1) and Young’s convolutional inequality.
@@
Proof of (C13). The weak compactness is the direct consequence of (E4) and the inequality $u_{\varepsilon}\leq s_{\varepsilon}$. The identity ${s^{l-m}}\overline{\,\frac{u^{k}}{s^{l}}\,}=\overline{\,\frac{u^{k}}{s^{m}}\,}$ follows by writing $s_{\varepsilon}^{l-m}\,\frac{u_{\varepsilon}^{k}}{s_{\varepsilon}^{l}}=\frac{u_{\varepsilon}^{k}}{s_{\varepsilon}^{m}}$ the limit, let $f_{\varepsilon}=\frac{u_{\varepsilon}^{k}}{s_{\varepsilon}^{l}}$ and passing to the limit in this identity using (C1), (C8) and the weak compactness proved in the current step.
Proof of (C14). The three convergences follow by the Banach-Alaoglu theorem and the bounds (E2) and (E4). For the first identity, we write $\frac{u_{\varepsilon}}{s_{\varepsilon}}\,\partial_{x}s_{\varepsilon}^{\alpha}+\frac{v_{\varepsilon}}{s_{\varepsilon}}\,\partial_{x}s_{\varepsilon}^{\alpha}=\partial_{x}s_{\varepsilon}^{\alpha}$ (which is immediate when $s_{\varepsilon}>0$ while when $s_{\varepsilon}=0$ we use that $\partial_{x}s_{\varepsilon}^{\alpha}=2\,s_{\varepsilon}^{\frac{\alpha}{2}}\,\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}=0$ since both derivatives are $L^{2}((0,T)\times\mathbb{R})$ functions) and we pass to the limit in the identity using (C3). Similarly, for the second identity we write $s_{\varepsilon}\,\frac{u_{\varepsilon}}{s_{\varepsilon}}\,\partial_{x}s_{\varepsilon}^{\alpha}=u_{\varepsilon}\,\partial_{x}s_{\varepsilon}^{\alpha}$ and we pass to the limit. Both sides converge in $L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ using (C1) with $p=2$ so we deduce $s\,\overline{\frac{u}{s}\,\partial_{x}s^{\alpha}}=\overline{u\,\partial_{x}s^{\alpha}}$. ∎
@@ 3.3. A conservation law applicable for all $\alpha\in(0,\infty)$
The main novel point in this section is the observation that there is a nonlinear function of $u_{\varepsilon}$ which satisfies a certain conservation law, suitable for application of our variant of the div-curl lemma (Theorem 4.3).
@@
$(\partial_{x}u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]+u_{\varepsilon}\,\partial_{x}^{2}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}])\,(1+\log u_{\varepsilon})=$
@@ 3.4. A conservation law applicable for $\alpha\in(0,2]$
Here, we provide another conservation law which is applicable only for $\alpha\in(0,2]$ (this range might be not optimal but certainly it cannot be $(0,\infty)$ as explained below). This conservation law can be used instead of (3.17) to prove Proposition 5.2 (and consequently, Theorem 1.3) slightly easier and we present this alternative argument in Section 5.4. The quantity that turns out to be useful for us, $\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}$, can be thought of as an integrand of the relative $L^{2}$ entropy $(1-\frac{u_{\varepsilon}}{s_{\varepsilon}})^{2}\,s_{\varepsilon}$ up to the terms whose limit can be identified. Note also that a similar relationship holds for $\frac{u_{\varepsilon}^{k+1}}{s_{\varepsilon}^{k}}$.
@@
Let $u_{\varepsilon}$, $v_{\varepsilon}$ solve (1.3) and let $s_{\varepsilon}=u_{\varepsilon}+v_{\varepsilon}$. Then, $\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}$ satisfies
@@
Moreover, $h_{\varepsilon}\to 0$ strongly in $L^{2}(0,T;H^{-1}(\mathbb{R}))$ and, if $\alpha\leq 2$, the sequence $\{g_{\varepsilon}\}_{\varepsilon\in(0,1)}$ is bounded in $L^{1}(0,T;L^{1}_{\operatorname{loc}}(\mathbb{R}))$.
We remark that the restriction $\alpha\leq 2$ is only necessary to guarantee that $\{g_{\varepsilon}\}_{\varepsilon\in(0,1)}$ is bounded in $L^{1}(0,T;L^{1}_{\operatorname{loc}}(\mathbb{R}))$ which is achieved by proving that $\{\partial_{x}s_{\varepsilon}\}_{\varepsilon\in(0,1)}$ is bounded in $L^{1}(0,T;L^{1}_{\operatorname{loc}}(\mathbb{R}))$. If one established this bound for $\alpha>2$, the conclusion of Proposition 3.7 would be also true in this case. Furthermore, the proof of Proposition 5.2 presented in Section 5.4 would also work, leading to a different (and simpler) proof of Theorem 1.3.
@@
$\begin{split}&\partial_{t}\left(\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\right)=\frac{2\,u_{\varepsilon}\,\partial_{t}u_{\varepsilon}\,s_{\varepsilon}-u_{\varepsilon}^{2}\,\partial_{t}s_{\varepsilon}}{s_{\varepsilon}^{2}}=\\ &=\,\frac{2\,u_{\varepsilon}\,s_{\varepsilon}}{s_{\varepsilon}^{2}}\,\left[\partial_{x}u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})+u_{\varepsilon}\,\partial_{x}^{2}p(s_{\varepsilon})+\partial_{x}(u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}])+\varepsilon\,\partial_{x}^{2}u_{\varepsilon}\right]\\ &-\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\left[\partial_{x}s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})+s_{\varepsilon}\,\partial_{x}^{2}p(s_{\varepsilon})+\partial_{x}(u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}])+\partial_{x}(v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])+\varepsilon\,\partial_{x}^{2}s_{\varepsilon}\right].\end{split}$ (3.20)

### 16 Apr 2026 v2

@@ 3.1. A priori estimates
$\|s_{\varepsilon}^{\frac{\gamma+\alpha}{2}+1}\|_{L^{\infty}(\mathbb{R})}\leq\|\partial_{x}s_{\varepsilon}^{\frac{\gamma+\alpha}{2}+1}\|_{L^{1}(\mathbb{R})}\leq\left(\frac{\gamma+\alpha}{2}+1\right)\|s_{\varepsilon}^{\frac{3}{2}}\|_{L^{2}(\mathbb{R})}\,\|s_{\varepsilon}^{\frac{\gamma+\alpha-3}{2}}\,\partial_{x}s_{\varepsilon}\|_{L^{2}(\mathbb{R})},$ (3.12)
@@
$\int_{\mathbb{R}}s_{\varepsilon}^{\frac{\gamma-3\alpha}{2}}\mathop{}\!\mathrm{d}x\leq\left(\int_{\mathbb{R}}s_{\varepsilon}^{\frac{\gamma}{2}}\mathop{}\!\mathrm{d}x\right)^{\beta}\,M^{1-\beta},\qquad\qquad\beta=\frac{\frac{\gamma-3\alpha}{2}-1}{\frac{\gamma}{2}-1},$ (3.14)
@@
We finally observe that $M^{2(1-\beta)}\leq 1+M^{2}$ and $\gamma\geq 1$ and $\beta\geq\frac{4}{5}$ imply $\gamma^{\frac{4}{\beta}}\leq\gamma^{5}$ so that
$\partial_{t}\int_{\mathbb{R}}s_{\varepsilon}^{\gamma}\mathop{}\!\mathrm{d}x\leq\frac{1+M^{2}}{\delta_{1}^{2}}\,\|s_{\varepsilon}^{\frac{3}{2}}\|^{2}_{L^{2}(\mathbb{R})}\,\left(1+\gamma^{5}\,\left(\int_{\mathbb{R}}s_{\varepsilon}^{\frac{\gamma}{2}}\mathop{}\!\mathrm{d}x\right)^{2}\right).$ (3.15)
@@
Step 3. Proof of (E5), (E6). Clearly, $\{u_{\varepsilon}\log u_{\varepsilon}\}$, $\{v_{\varepsilon}\log v_{\varepsilon}\}$ are bounded in $L^{\infty}((0,T)\times\mathbb{R})$ by (E4). Hence, to obtain (E5) we need to prove that $\{u_{\varepsilon}\log u_{\varepsilon}\}$, $\{v_{\varepsilon}\log v_{\varepsilon}\}$ are bounded in $L^{\infty}(0,T;L^{1}(\mathbb{R}))$. We observe that by Lemma 3.3 with $|y|^{-}=-\min(0,y)$
$\int_{\mathbb{R}}u_{\varepsilon}|\log u_{\varepsilon}|\mathop{}\!\mathrm{d}x=\int_{\mathbb{R}}u_{\varepsilon}\log u_{\varepsilon}\mathop{}\!\mathrm{d}x+2\int_{\mathbb{R}}u_{\varepsilon}|\log u_{\varepsilon}|^{-}\mathop{}\!\mathrm{d}x\leq\\ \leq\int_{\mathbb{R}}u_{\varepsilon}\log u_{\varepsilon}\mathop{}\!\mathrm{d}x+2\left(\int_{\mathbb{R}}u_{\varepsilon}\,|x|\mathop{}\!\mathrm{d}x+\frac{2}{e}\int_{\mathbb{R}}e^{-|x|}\mathop{}\!\mathrm{d}x\right).$
The same inequality holds for $v_{\varepsilon}$. Therefore, the claim follows from (3.2) and (E1).
To prove (E6), we split the integral for three sets as follows:
$\int_{\mathbb{R}}u_{\varepsilon}|\log u_{\varepsilon}||x|^{\frac{1}{2}}\mathop{}\!\mathrm{d}x\leq\int_{u_{\varepsilon}\geq 1}u_{\varepsilon}|\log u_{\varepsilon}||x|^{\frac{1}{2}}\mathop{}\!\mathrm{d}x+\int_{u_{\varepsilon}\in[e^{-|x|^{1/2}},1]}u_{\varepsilon}|\log u_{\varepsilon}||x|^{\frac{1}{2}}\mathop{}\!\mathrm{d}x\\ +\int_{u_{\varepsilon}\leq e^{-|x|^{1/2}}}u_{\varepsilon}|\log u_{\varepsilon}||x|^{\frac{1}{2}}\mathop{}\!\mathrm{d}x=:I_{1}+I_{2}+I_{3}.$
For $I_{1}$, we can estimate $|\log u_{\varepsilon}|\leq 1+u_{\varepsilon}\leq 1+\|u_{\varepsilon}\|_{L^{\infty}}$ by (E4) and $|x|^{\frac{1}{2}}\leq 1+|x|$ so we can use (E1) to control the integral. For $I_{2}$, the set of integration implies $|\log u_{\varepsilon}|\leq|x|^{\frac{1}{2}}$ so we can use again (E1) to control the term. Finally, to bound $I_{3}$, we write $u_{\varepsilon}|\log u_{\varepsilon}||x|^{\frac{1}{2}}=\sqrt{u_{\varepsilon}}\,\sqrt{u_{\varepsilon}}|\log u_{\varepsilon}|\,|x|^{\frac{1}{2}}\leq C\,|x|^{\frac{1}{2}}\,e^{-|x|^{\frac{1}{2}}/2}$ where $C=\sup_{x\in[0,1]}\sqrt{x}\,|\log x|$. The proof for $v_{\varepsilon}$ is the same.
Step 4. Proof of (E3) and (E8). For (E3), we observe that
@@
Step 5. Proof of (E9). We have $\partial_{t}u_{\varepsilon}=\partial_{x}F_{\varepsilon}$ where
@@
Step 6. Proof of (E10). Since $\alpha\leq 2$ we use that $\partial_{x}s_{\varepsilon}=\frac{2}{\alpha}\,s_{\varepsilon}^{1-\frac{\alpha}{2}}\,\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}$ so $\{\partial_{x}s_{\varepsilon}\}$ is bounded in $L^{2}((0,T)\times\mathbb{R})$ by (E2) and (E4). ∎
@@ 3.2. Basic compactness
    $\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},\!v_{\varepsilon}]\!\to\!\partial_{x}\mathcal{V}^{1}[u,\!v],\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},\!v_{\varepsilon}]\!\to\!\partial_{x}\mathcal{V}^{2}[u,\!v]$ a.e. and strongly in $L^{p}_{\operatorname{loc}}((0,T)\!\times\!\mathbb{R})$ for $p\in[1,\infty)$,
@@
Concerning the overline notation, we use the it to denote weak limits which exist by the a priori estimates but we do not know their representation. In particular, we do not know that $\overline{|\partial_{x}s^{\alpha}|^{2}}=|\partial_{x}s^{\alpha}|^{2}$ or $\overline{u\,\partial_{x}s^{\alpha}}=u\,\partial_{x}s^{\alpha}$. Before proceeding to the proof, we will state a version of the Aubin-Lions Lemma taken from [52, Corollary 7.9] in the case that the time derivative is a measure.
@@
$\frac{1}{1+\frac{\alpha}{2}}\int_{0}^{T}\int_{\mathbb{R}}\partial_{t}s_{\varepsilon}^{1+\frac{\alpha}{2}}\,\varphi\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}t\leq\|\partial_{t}s_{\varepsilon}\|_{L^{2}(0,T;H^{-1}(\mathbb{R}))}\,\|s_{\varepsilon}^{\frac{\alpha}{2}}\|_{L^{2}(0,T;H^{1}(\mathbb{R}))}\,\|\varphi\|_{C(0,T;W^{1,\infty}_{0}(B_{R}))},$
@@
$\|\partial_{x}K^{1,1}\ast(u_{\varepsilon}-u)\|_{L^{1}((0,T)\times\Omega)}\leq\\ \leq\|\partial_{x}K^{1,1}\ast((u_{\varepsilon}-u)\,\chi_{R})\|_{L^{1}((0,T)\times\Omega)}+\|\partial_{x}K^{1,1}\ast((u_{\varepsilon}-u)\,(1-\chi_{R}))\|_{L^{1}((0,T)\times\Omega)}.$
The first term converges to 0 for each $R>0$ (indeed, by a standard version of the Aubin-Lions lemma, it has a subsequence converging to 0 so we deduce that the whole sequence converges to 0; in order to apply the Aubin-Lions lemma, the spatial regularity is obtained from $\partial_{x}K^{1,1},\partial_{x}^{2}K^{1,1}\in L^{\infty}(\mathbb{R})$, the temporal regularity is guaranteed by (E9) and the whole reasoning is restricted to a compact set thanks to the function $\chi_{R}$ and $\Omega$ being bounded). The second term is arbitrarily small by the assumption $\partial_{x}K^{1,1}\in L^{\infty}(\mathbb{R})$, the moment estimate (E1) and Young’s convolutional inequality. As soon as the convergence in $L^{1}((0,T)\times\Omega)$ is established, we can extend it to $L^{p}((0,T)\times\Omega)$ with $1<p<\infty$ by interpolation while the a.e. convergence follows by extracting a further subsequence.
@@
Proof of (C13). The weak compactness is the direct consequence of (E4) and the inequality $u_{\varepsilon}\leq s_{\varepsilon}$. The identity ${s^{l-m}}\overline{\,\frac{u^{k}}{s^{l}}\,}=\overline{\,\frac{u^{k}}{s^{m}}\,}$ follows by writing $s_{\varepsilon}^{l-m}\,\frac{u_{\varepsilon}^{k}}{s_{\varepsilon}^{l}}=\frac{u_{\varepsilon}^{k}}{s_{\varepsilon}^{m}}$ the limit, let $f_{\varepsilon}=\frac{u_{\varepsilon}^{k}}{s_{\varepsilon}^{l}}$ and passing to the limit in this identity using (C1) and the weak compactness proved in the current step.
Proof of (C14). The four convergences follow by the Banach-Alaoglu theorem and the bounds (E3) and (E4). For the first identity, we write $\frac{u_{\varepsilon}}{s_{\varepsilon}}\,\partial_{x}s_{\varepsilon}^{\alpha}+\frac{v_{\varepsilon}}{s_{\varepsilon}}\,\partial_{x}s_{\varepsilon}^{\alpha}=\partial_{x}s_{\varepsilon}^{\alpha}$ (which is immediate when $s_{\varepsilon}>0$ while when $s_{\varepsilon}=0$ we use that $\partial_{x}s_{\varepsilon}^{\alpha}=2\,s_{\varepsilon}^{\frac{\alpha}{2}}\,\partial_{x}s_{\varepsilon}^{\frac{\alpha}{2}}=0$ since both derivatives are $L^{2}((0,T)\times\mathbb{R})$ functions) and we pass to the limit in the identity using (C3). Similarly, for the second identity we write $s_{\varepsilon}\,\frac{u_{\varepsilon}}{s_{\varepsilon}}\,\partial_{x}s_{\varepsilon}^{\alpha}=u_{\varepsilon}\,\partial_{x}s_{\varepsilon}^{\alpha}$ and we pass to the limit. Both sides converge in $L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ using (C1) with $p=2$ so we deduce $s\,\overline{\frac{u}{s}\,\partial_{x}s^{\alpha}}=\overline{u\,\partial_{x}s^{\alpha}}$. ∎
@@ 3.3. A conservation law applicable for all $\alpha\in(0,\infty)$
The main novel point in this section is the observation that there is a nonlinear function of $u_{\varepsilon}$ which satisfies a certain conservation law, suitable for application of the Div-Curl Lemma (Theorem 4.3).
@@
$(\partial_{x}u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]+u_{\varepsilon}\,\partial_{x}^{2}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}])\,(1+\log u_{\varepsilon})$
@@ 3.4. A conservation law applicable for $\alpha\in(0,2]$
Here, we provide another conservation law which is applicable only for $\alpha\in(0,2]$ (see Remark 3.8 for a discussion of the range). This conservation law can be used instead of (3.17) to prove Proposition 5.2 (and consequently, Theorem 1.3) slightly easier and we present this alternative argument in Section 5.4. The quantity that turns out to be useful for us, $\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}$, can be thought of as an integrand of the relative $L^{2}$ entropy $(1-\frac{u_{\varepsilon}}{s_{\varepsilon}})^{2}\,s_{\varepsilon}$ up to the terms whose limit can be identified. Note also that a similar relationship holds for $\frac{u_{\varepsilon}^{k+1}}{s_{\varepsilon}^{k}}$ (see Remark 3.9 below).
@@
Let $u_{\varepsilon}$, $v_{\varepsilon}$ solve (1.4) and let $s_{\varepsilon}=u_{\varepsilon}+v_{\varepsilon}$. Then,
@@
Moreover, $h_{\varepsilon}\to 0$ strongly in $L^{2}(0,T;H^{-1}(\mathbb{R}))$ and, if $\alpha\leq 2$, the sequence $\{g_{\varepsilon}\}_{\varepsilon\in(0,1)}$ is bounded in $L^{1}(0,T;L^{1}_{\operatorname{loc}}(\mathbb{R}))$ (see Remark 3.8 below for a discussion of the condition $\alpha\leq 2$).
@@
$\begin{split}&\partial_{t}\left(\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\right)=\frac{2\,u_{\varepsilon}\,\partial_{t}u_{\varepsilon}\,s_{\varepsilon}-u_{\varepsilon}^{2}\,\partial_{t}s_{\varepsilon}}{s_{\varepsilon}^{2}}\\ &=\,\frac{2\,u_{\varepsilon}\,s_{\varepsilon}}{s_{\varepsilon}^{2}}\,\left[\partial_{x}u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})+u_{\varepsilon}\,\partial_{x}^{2}p(s_{\varepsilon})+\partial_{x}(u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}])+\varepsilon\,\partial_{x}^{2}u_{\varepsilon}\right]\\ &\phantom{=}\,-\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\left[\partial_{x}s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})+s_{\varepsilon}\,\partial_{x}^{2}p(s_{\varepsilon})+\partial_{x}(u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}])+\partial_{x}(v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])+\varepsilon\,\partial_{x}^{2}s_{\varepsilon}\right].\end{split}$ (3.20)